OpenStudy (anonymous):

I'm using the chain rule and i need to find the derivative of Y=((3x-2)/(6-5x))^4

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

do you know how to find the derivative of \[\frac{3x-2}{6-5x}\]?

OpenStudy (anonymous):

that is all the hard work.

OpenStudy (anonymous):

then the "chain rule" will tell you that since this is something to the power of 4, its derivative will be 4 something cubed times the derivative of something. so all the work is finding the derivative of that quotient.

OpenStudy (anonymous):

i can write it for you if you like

OpenStudy (anonymous):

hello anwar. feel like writing it out?

OpenStudy (anonymous):

\[{d \over dx}({3x-2 \over (6-5x)^4})={3(6-5x)-4(3x-2)(6-5x)^3(-5) \over (6-5x)^8}\] Simplify!!

OpenStudy (anonymous):

hello satellite :)

OpenStudy (anonymous):

ok i will. i get \[\frac{8}{(5x-6)^2}\]

OpenStudy (anonymous):

The denominator is raised to the power of 4, isn't it?

OpenStudy (anonymous):

so back to the problem. the derivative will be \[4(\text{inside thing})^3\times \frac{8}{(5x-6)^2}\]

OpenStudy (anonymous):

Oh the whole thing.. my bad.

OpenStudy (anonymous):

no i was assuming it was the whole thing raised to the 4th. chain rule and all that.

OpenStudy (anonymous):

never mind my solution doshi, just stick with satellite :D

OpenStudy (anonymous):

inside piece just \[\frac{3x-2}{5x-6}\]

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

well actually it is \[\frac{3x-2}{6-5x}\]

OpenStudy (anonymous):

whose derivative is \[\frac{3(6-5x)+5(3x-2)}{(5x-6)^2}=\frac{8}{(5x-6)^2}\]

OpenStudy (anonymous):

The derivative \[y'=4({3x-2 \over 6-5x})^3{8 \over (6-5x)^2}\]

OpenStudy (anonymous):

bet we lost doshi somewhere

OpenStudy (anonymous):

probably gave up in disgust

OpenStudy (anonymous):

I think so :D